If Veronica Vaughn spends all of her daily income on cigarettes and Yoo Hoo, she can afford 10 packs of cigarettes and 10 bottles of Yoo Hoo. She can also afford 6 packs of cigarettes and 22 bottles of Yoo Hoo.


Now suppose the price of cigarettes falls by $1 and the price of YooHoo roses by $1. If Veronica was consuming 5 packs of cigarettes and 30 bottles of Yoo Hoo prior to the price changes, how much must her income rise under the new prices in order for her to just afford the old bundle, (5,30)?


1
Expert's answer
2021-11-09T10:45:45-0500

Solution:

M = PxX + PyY

Where: M = Income

            Px = Price of cigarettes

            Py = Price of Yoo Hoo

             X = Cigarettes

             Y = Yoo Hoo

M = 10Px + 10Py

M = 6Px + 22Py

M/10 = Px

M/10 = Py

Assume M = 120

Px = "\frac{120}{10} = 12"


Py = "\frac{120}{10} = 12"

 

New price of cigarettes = 12 – 1 = 11

New price of Yoo Hoo = 12 + 1 = 13

 

New income = (5 "\times" 11) + (30 "\times" 13) = 55 + 390 = 445